# Enhancing LLM Reasoning via Non-Human-Like Reasoning Path Preference Optimization

### v1

@@ 4 Experiments and Analysis
In this section, we first introduce our experimental setup, including datasets, models, baselines, metrics, and parameter settings. We then present the experimental results, followed by an analysis of the CGPO transferability, generalization, and scalability.
@@ 4.1 Experiments Setup
Datasets and Models: For mathematical reasoning tasks, we use the prompts from the Step-DPO-10k [^1] https://huggingface.co/datasets/xinlai/Math-Step-DPO-10K dataset as model inputs $\mathbf{x}$ to construct the preference dataset for CGPO training. MetaMath-Mistral-7B (Yu et al., 2024), MetaMath-LLaMA-8B (Liu et al., 2025), Qwen2-7B-SFT [^2] https://huggingface.co/xinlai/Qwen2-7B-SFT, Qwen1.5-32B-SFT [^3] https://huggingface.co/xinlai/Qwen1.5-32B-SFT and DeepSeekMath-Base-SFT [^4] https://huggingface.co/xinlai/DeepSeekMath-Base-SFT are employed as the policy models $\pi_{0}$. To ensure that the capabilities of the Reward Model (RM) align with the policy model and can provide token-level fine-grained reward, we adopt ASPRM-L [^5] https://huggingface.co/Lux0926/ASPRM-L as $R$ to select candidate tokens after $s_{init}$, preventing performance gains from an overly strong RM.
@@
For code generation tasks, we use the prompts from the training set of the LeetCodeDataset as model inputs $\mathbf{x}$ to construct the CGPO preference dataset. The policy model $\pi_{0}$ is Deepseek-Coder-7B-Instruct-v1.5 (Guo et al., 2024) and we adopt ASPRM-D [^6] https://huggingface.co/Lux0926/ASPRM-D as our reward model to select candidate tokens. The trained models are evaluated on the test sets of LiveCodeBench [^7] https://huggingface.co/livecodebench and LeetCodeDataset [^8] https://huggingface.co/datasets/newfacade/LeetCodeDataset.
Baselines and Metrics: For the mathematics domain, two baselines are considered: the base model and the model trained using the Step-DPO method. For the code domain, only the base model serves as the baseline.
@@
For all tasks, model performance is reported using the accuracy metric. In mathematical reasoning tasks, a prediction is considered correct only if model’s final answer exactly matches the ground truth. For code-generation tasks, the generated code is executed in a sandbox and deemed correct if it passes all test cases.
@@
Parameter Settings: For each base model in the mathematical domain, we construct the training data as follows. All 10,795 prompts from Step-DPO-10k are fed through the model to generate one output per prompt, yielding 10,795 outputs. For each output, we locate the lowest confidence point and use the reward model to score the top 8 token candidates, then branch into two reasoning paths. Decoding proceeds until the confidence falls below the 2% of the model’s confidence distribution or <EOS> token is reached. This produces a preference dataset of 10,795 pairs for each base model, equivalent in size to Step-DPO-10k. For the code domain base model, we sample 4 times per prompt from the 2,641 LeetCodeDataset training examples and apply the same token selection and reward-model procedure to generate 10,564 preference pairs.
For MetaMath-Mistral-7B and MetaMath-Mistral-8B, we follow the training configuration released by Step-DPO, set $\beta$ to 0.4, use a learning rate of 5e-7, set the global batch size to 128, and train for 8 epochs to reproduce the Step-DPO method. For all other Step-DPO baselines, we use the publicly released model checkpoints by Step-DPO. For all CGPO models, we train for 4 epochs using a global batch size of 128 and a learning rate of 5e-7. For MetaMath-Mistral-7B, Qwen2-7B-SFT and Deepseek-Coder-7B-Instruct-v1.5, we set $\beta$ to 0.4. For other models, we set $\beta$ to 0.3. We use the AdamW optimizer and a cosine learning rate schedule, with a warmup ratio of 0.1.
@@
Dataset Model Base + Step-DPO + CGPO
GSM8K MetaMath-Mistral-7B 77.1 76.7 (-0.4) 79.4 (+2.3)
MetaMath-Llama-8B 81.8 84.8 (+3.0) 85.3 (+3.5)
Qwen2-7B-SFT 87.9 88.4 (+0.5) 88.6 (+0.7)
Qwen1.5-32B-SFT 90.0 90.3 (+0.3) 90.4 (+0.4)
DeepSeekMath-Base-7B-SFT 86.3 86.5 (+0.2) 87.1 (+0.8)
MATH MetaMath-Mistral-7B 26.6 27.3 (+0.7) 27.0 (+0.4)
MetaMath-Llama-8B 39.3 40.2 (+0.9) 40.3 (+1.0)
Qwen2-7B-SFT 54.2 54.6 (+0.4) 54.8 (+0.6)
Qwen1.5-32B-SFT 54.2 55.7 (+1.5) 55.2 (+1.0)
DeepSeekMath-Base-7B-SFT 51.4 52.5 (+1.1) 51.8 (+0.4)
LiveCodeBench Deepseek-Coder-7B-Instruct-v1.5 19.3 / 19.7 (+0.4)
LeetCodeDataset Deepseek-Coder-7B-Instruct-v1.5 12.7 / 13.2 (+0.5)
@@ 4.2 Overall Results
The main results are shown in Table 1. For the mathematics reasoning task, our method outperforms Step-DPO on GSM8K and achieves the greatest improvement on MetaMath-Llama-8B, with a 4.3% increase compared to the base model. We attribute this phenomenon to the reward model and the policy model being trained on the same series of foundational models. Notably, when training MetaMath-Mistral-7B using the Step-DPO method, the resulting model’s performance actually declined compared to the base model. This may be because it is difficult for the weaker model to follow the human-like reasoning paths and learn from them in a preference way. However, our method still leads to an improvement compared to the base model, which demonstrates the advantage of the model learning from non-human-like reasoning paths through self-exploration. On the MATH dataset, our method outperforms Step-DPO on MetaMath-Llama-8B and Qwen2-7B-SFT, but performs less well on the other three models. That may be because the MATH dataset is more challenging than the GSM8K dataset. The 10k data volume is insufficient for fully exploring non-human-like reasoning paths, limiting the model improvement. Additional analysis experiments are conducted below to verify this.
Additionally, for the code generation task, the Deepseek-Coder-7B-Instruct-v1.5 trained with our method achieved 2.1% and 4.0% over the base model on the LiveCodeBench and LeetCodeDataset respectively. This demonstrates that, beyond mathematical reasoning tasks, CGPO can also enhance model capabilities in tasks such as code generation, where human-like reasoning paths are not easily defined, by enabling the model to explore non-human-like reasoning paths.
Table 2: Scalability on CGPO training data: models were trained with 10k, 20k, 40k, and 80k data samples, respectively.
Model Dataset Data Scale
10k 20k 40k 80k
MetaMath-Llama-8B GSM8K 85.3 85.3 85.6 86.4
MATH 40.3 40.8 41.2 41.2
Qwen2-7B-SFT GSM8K 88.6 88.7 88.9 89.5
MATH 54.8 54.9 55.1 55.3
@@ 4.3 Analysis of Scalability
Table 2 presents model performance on the GSM8K and MATH datasets. We find that increasing the exploration degree gradually improves performance on both datasets, with a notable jump on GSM8K when the data volume increases from 40K to 80K. These results indicate that CGPO exhibits effective scalability.
Table 3: Performance comparison on out-of-domain dataset (Omni-Math).
Dataset Model Base + Step-DPO + CGPO
Omni-Math MetaMath-Mistral-7B 8.4 7.7 (-0.7) 8.8 (+0.4)
MetaMath-Llama-8B 11.4 11.1 (-0.3) 11.7 (+0.3)
Qwen2-7B-SFT 14.3 14.0 (-0.3) 14.6 (+0.3)
DeepSeekMath-Base-7B-SFT 13.7 14.7 (+1.0) 14.2 (+0.5)
Qwen1.5-32B-SFT 14.5 15.1 (+0.6) 15.2 (+0.7)
4.4 Analysis of Out of Datasets
To analyze the out-of-distribution generalization performance of CGPO and its performance on challenging mathematical tasks, we use the Omni-Math dataset (Gao et al., 2024a) is used to evaluate the models. Omni-Math is an evaluation benchmark that encompasses 4,428 Olympic-level competition problems of 10 different difficulty levels and 33 subfields.
As shown in Table 3, compared with Step-DPO, CGPO performs slightly lower only on DeepSeekMath-Base-7B-SFT, while outperforms it on all other models. This shows that CGPO exhibits superior out-of-distribution generalization and can effectively enhance the performance of weaker models on challenging benchmarks, even without strong model annotations.
Table 4: Results of training datasets constructed by different reward models: ASPRM and Math-Shepherd denote the models trained with preference data guided by these two reward models.
@@ 4.4 Analysis of Out of Datasets
Qwen2-7B-SFT   + ASPRM 88.6 54.8
@@
4.5 Analysis of Reward Model
To explore the impact of different reward models on CGPO, we adopt the process reward model proposed by Math-Shepherd (Wang et al., 2024a) for comparison with ASPRM. The comparison results are presented in Table 4.
When Math-Shepherd is used instead of ASPRM, the performance of CGPO-trained model declines, although it remains better than the base model. Even with a weaker PRM, CGPO still shows an improvement over the base model. This improvement stems from precise token-level rewards, underscoring the significance of fine-grained evaluation.
@@ 4.5 Analysis of Reward Model
4.6 Analysis of Stopping Threshold
We report the performance differences of the model under different confidence thresholds at which the sampling stops in Figure 3. For MetaMath-Llama-8B, performance keeps improving as the threshold increases. For Qwen2-7B-SFT, overall performance also shows an upward trend as the threshold rises, but the peak is achieved at a confidence threshold of 8%.
When we increase the confidence threshold, more samples will stop in the intermediate process rather than the generation of the <EOS> tag. This increases the proportion of non-human-like data, thereby enhancing the model’s exploration and learning of non-human-like reasoning paths, which ultimately improves performance. However, increasing the confidence threshold also introduces shorter lengths of chosen and rejected data, as shown in Figure 4. We argue that this is the reason for the performance decline of Qwen2-7B-SFT when the confidence threshold is greater. These results further demonstrate the necessity of optimizing the model by the intermediate reasoning path rather than by the entire process.
4.7 Analysis of Division Strategy
We use confidence scores to reconstruct the initial reasoning step , as well as the chosen and rejected data in the original Step-DPO-10k dataset, in order to compare human-like and non-human-like step divisions. Specifically, we truncate the initial reasoning step in the Step-DPO-10k dataset at the lowest confidence point, as in the dataset construction of CGPO. For the chosen and rejected data, we apply a truncation threshold at the lowest 2% confidence. The final results are reported in Table 5.
We find that even using $\pi_{0}$ to split the reasoning path on outputs generated by the strong model for training, the resulting performance still surpasses the baseline. However, it does not outperform training on the $\pi_{0}$ own generated and split data. This result suggests that the split of its own generated outputs by the weak model is more aligned with the model, thereby leading to improved training performance.
Model GSM8K MATH
MetaMath-Llama-8B   + CGPO 85.3 40.3
+ Step-DPO 84.8 40.2
+ Step-DPO-Confidence 84.4 40.2
Qwen2-7B-SFT   + CGPO 88.6 54.8
+ Step-DPO 88.4 54.6
+ Step-DPO-Confidence 88.1 54.5

### v2

@@ 4 Experiments and Analysis
In this section, we first introduce our experimental setup, including datasets, models, baselines, metrics, and parameter settings. We then present the experimental results, followed by an analysis of the transferability, generalization, and scalability of CGPO.
@@ 4.1 Experiments Setup
Datasets and Models: For mathematical reasoning tasks, we use the prompts from the Step-DPO-10k [^1] https://huggingface.co/datasets/xinlai/Math-Step-DPO-10K dataset as the model inputs $\mathbf{x}$ to construct the preference dataset for training. We evaluate CGPO with various models serving as the policy model $\pi_{\theta}$, including MetaMath-Mistral-7B (Yu et al., 2024), MetaMath-Llama-8B (Liu et al., 2025), Qwen2-7B-SFT, Qwen1.5-32B-SFT, DeepSeekMath-Base-SFT(Lai et al., 2024), DeepSeek-R1-Distill-Qwen-1.5B, and DeepSeek-R1-Distill-Qwen-7B(Guo et al., 2025). To ensure that the capabilities of the reward model (RM) $R$ align with the policy model to provide accurate token-level reward, to prevent performance gains from an overly strong RM, We adopt ASPRM-L [^2] https://huggingface.co/Lux0926/ASPRM-L, whose base model is competitive with our policy model while enabling fine-grained evaluation, as the RM to select candidate tokens after $\textbf{s}_{init}$. For evaluation, we benchmark standard instruction models on the GSM8K and MATH test sets, while reasoning models are evaluated on GSM8K and MATH500 to enhance evaluation efficiency.
@@
For code generation tasks, we use the prompts from the training set of the LeetCodeDataset [^3] https://huggingface.co/datasets/newfacade/LeetCodeDataset as model inputs $\mathbf{x}$ to construct the preference dataset. The policy model $\pi_{\theta}$ is Deepseek-Coder-7B-Instruct-v1.5 (Guo et al., 2024), and we adopt ASPRM-D [^4] https://huggingface.co/Lux0926/ASPRM-D as our reward model to select candidate tokens. The trained models are evaluated on the test sets of LiveCodeBench-v6555https://huggingface.co/livecodebench and LeetCodeDataset.
Baselines and Metrics: For the mathematics domain, five training settings are compared: the base model, Step-DPO, CGPO, DPO trained on the sampled full chosen and rejected sequence pairs constructed by Step-DPO (Step-DPO data), and DPO trained on the sampled full chosen and rejected sequence pairs constructed by CGPO (CGPO data). For the code domain, three training settings are compared: the base model, DPO trained on CGPO data, and CGPO.
@@
For all tasks, large reasoning models follow prior work(Hugging Face, 2025) by sampling 4 generations with temperature $0.6$ and top-$p$ $0.95$, and reporting pass@1 for comparison; all other models use greedy decoding and are compared by directly computed accuracy. In mathematical reasoning tasks, a prediction is considered correct only if model’s final answer exactly matches the ground truth. For code-generation tasks, the generated code is executed in a sandbox and deemed correct if it passes all test cases.
@@
Parameter Settings: For each base model in the mathematical domain, we sample one time per prompt from the dataset, yielding 10,795 outputs. For each output, we locate the lowest confidence point and use the reward model to score the top 8 token candidates, then branch into two reasoning paths. The policy models decode tokens until the confidence falls below 2% of its confidence distribution or reach EOS token. This produces a preference dataset of 10,795 pairs for each policy model, equivalent in size to Step-DPO-10k. For the code domain, we sample 4 times per prompt from the 2,641 LeetCodeDataset training data and apply the same token selection and reward-model procedure to generate 10,564 preference pairs. We follow the Step-DPO training configuration to reproduce baselines. Hyperparameter settings for CGPO are provided in Appendix A.3.
@@
Dataset Model Base + DPO + Step-DPO + CGPO
Step-DPO DATA CGPO DATA
GSM8K MetaMath-Mistral-7B 77.1 76.5 (-0.6) 78.5 (+1.4) 76.7 (-0.4) 81.1 (+4.0)
MetaMath-Llama-8B 81.8 85.0 (+3.2) 86.0 (+4.2) 84.8 (+3.0) 86.4 (+4.6)
Qwen2-7B-SFT 87.9 85.0 (-2.9) 88.8 (+0.9) 88.4 (+0.5) 89.8 (+1.9)
Qwen1.5-32B-SFT 90.0 90.3 (+0.3) 90.1 (+0.1) 90.3 (+0.3) 90.6 (+0.6)
DeepSeekMath-Base-7B-SFT 86.3 87.0 (+0.7) 86.9 (+0.6) 86.5 (+0.2) 87.4 (+1.1)
DeepSeek-R1-Distill-Qwen-1.5B 76.7 77.2 (+0.5) 76.3 (-0.4) 78.6 (+1.9) 79.0 (+2.3)
DeepSeek-R1-Distill-Qwen-7B 86.1 86.7 (+0.6) 85.6 (-0.5) 86.2 (+0.1) 86.5 (+0.4)
Average 83.7 84.0 (+0.3) 84.6 (+0.9) 84.5 (+0.8) 85.8 (+2.1)
MATH MetaMath-Mistral-7B 26.6 26.9 (+0.3) 27.0 (+0.4) 27.3 (+0.7) 27.4 (+0.8)
MetaMath-Llama-8B 39.3 40.1 (+0.8) 40.6 (+1.3) 40.2 (+0.7) 40.7 (+1.4)
Qwen2-7B-SFT 54.2 51.7 (-2.5) 54.5 (+0.3) 54.6 (+0.4) 55.0 (+0.8)
Qwen1.5-32B-SFT 54.2 54.4 (+0.2) 55.3 (+1.1) 55.5 (+1.3) 55.7 (+1.5)
DeepSeekMath-Base-7B-SFT 51.4 51.8 (+0.4) 51.9 (+0.5) 52.5 (+1.1) 52.8 (+1.4)
DeepSeek-R1-Distill-Qwen-1.5B 82.2 82.5 (+0.3) 82.3 (+0.1) 81.7 (-0.5) 83.7 (+1.5)
DeepSeek-R1-Distill-Qwen-7B 93.2 93.5 (+0.3) 93.3 (+0.1) 93.3 (+0.1) 94.0 (+0.8)
Average 57.3 57.3 (+0.0) 57.8 (+0.5) 57.9 (+0.6) 58.5 (+1.2)
LiveCodeBench Deepseek-Coder-7B-Instruct-v1.5 19.3 / 19.6 (+0.3) / 20.0 (+0.7)
LeetCodeDataset Deepseek-Coder-7B-Instruct-v1.5 12.7 / 12.8 (+0.1) / 13.6 (+0.9)
@@ 4.2 Overall Results
We show the main results in Table 1 and make the following observations: 1) For the mathematics reasoning task, CGPO achieves the highest average performance improvement across all models. Specifically, on GSM8K, CGPO yields the strongest performance for all evaluated base models except DeepSeek-R1-Distill-Qwen-7B, while on MATH, CGPO consistently achieves the best performance across all models. Compared to the base model, CGPO improves the average accuracy by 2.1% on GSM8K and 1.2% on MATH, indicating stable gains across different model families. Notably, Step-DPO leads to performance drops when applied to weaker models such as MetaMath-Mistral-7B and DeepSeek-R1-Distill-Qwen-1.5B, suggesting that alignment to human-like reasoning paths imposes a higher capability requirement and may yield suboptimal preference signals when the model struggles to follow such trajectories. 2) In addition, directly training DPO on full trajectories sampled by Step-DPO or CGPO can also degrade performance for certain models, possibly because sequence-level preference learning diffuses the supervision signal over the entire trajectory and fails to emphasize the critical reasoning bottlenecks. In contrast, CGPO produces consistent improvements over the base model across all evaluated settings, supporting the hypothesis that allowing models to explore and learn from non-human-like reasoning paths provides a more effective and optimizable search space than human-like reasoning paths. 3) Moreover, DPO trained on CGPO data generally outperforms its counterpart trained on Step-DPO data and, in several cases, approaches the gains of Step-DPO, further indicating that non-human-like, confidence-guided branching yields higher-quality preference data than alignment to human-like reasoning paths.
Additionally, for the code generation task, DeepSeek-Coder-7B-Instruct-v1.5 trained with CGPO achieved 0.7% improvement on LiveCodeBench and 0.9% improvement on the LeetCodeDataset. This demonstrates that CGPO can enhance model capabilities beyond mathematical reasoning tasks, including code generation. In such tasks, where human-like reasoning paths are not easily defined, CGPO enables the model to explore non-human-like reasoning paths and improves performance.
Table 2: Performance comparison on out-of-domain dataset (CNMO24).
Dataset Model Base + DPO + Step-DPO + CGPO
Step-DPO DATA CGPO DATA
CNMO24 MetaMath-Mistral-7B 0.0 0.0 (+0.0) 0.0 (+0.0) 0.0 (+0.0) 2.8 (+2.8)
MetaMath-Llama-8B 0.0 0.0 (+0.0) 0.0 (+0.0) 0.0 (+0.0) 2.8 (+2.8)
Qwen2-7B-SFT 5.6 0 (-5.6) 5.6 (+0.0) 5.6 (+0.0) 11.1 (+5.5)
DeepSeekMath-Base-7B-SFT 5.6 0 (-5.6) 5.6 (+0.0) 5.6 (+0.0) 5.6 (+0.0)
Qwen1.5-32B-SFT 5.6 2.8 (-2.8) 5.6 (+0.0) 8.3 +(2.7) 8.3 +(2.7)
DeepSeek-R1-Distill-Qwen-1.5B 8.3 8.3 (+0.0) 8.3 (+0.0) 5.6 (-2.7) 11.1 (+2.8)
DeepSeek-R1-Distill-Qwen-7B 8.3 8.3 (+0.0) 8.3 (+0.0) 5.6 (-2.7) 11.1 (+2.8)
Average 4.8 2.8 (-2.0) 4.8 (+0.0) 4.4 (-0.4) 7.5 (+2.8)
4.3 Analysis of Out-of-Distribution Datasets
To analyze the out-of-distribution generalization performance of CGPO and its performance on challenging multilingual mathematical tasks, the CNMO24 (Liu et al., 2024) dataset is used to evaluate the models. CNMO24 is a competition-level benchmark consisting of 18 problems from the China National Mathematical Olympiad (CNMO) 2024. We evaluate models on both the Chinese and English versions of each problem, yielding 36 evaluation instances in total.
As shown in Table 2, CGPO achieves the largest average improvement across all evaluated models compared with other settings. In particular, for weaker models such as MetaMath-Mistral-7B and MetaMath-Llama-8B, CGPO improves performance from 0% to a non-zero success rate. This shows that CGPO exhibits superior out-of-distribution generalization and can effectively enhance the performance of weaker models on challenging benchmarks, even without strong model guidance.
@@ 4.4 Analysis of Scalability
Figure 3: The relationship between the training dataset scale and the trained model performance.
@@
Figure 3 presents model performance on the GSM8K and MATH datasets. We find that increasing the exploration improves performance on both datasets, with a notable jump on GSM8K when the data volume increases from 40K to 80K. This suggests that increasing the training data enables models to explore more non-human-like reasoning paths, which further improves performance. These results indicate that CGPO exhibits effective scalability.
4.5 Analysis of Reward Model
To explore the impact of different reward models on CGPO, we adopt the process reward model proposed by Math-Shepherd (Wang et al., 2024a) and Qwen-2.5-Math-PRM(Zhang et al., 2025b) for comparison with ASPRM. To completely eliminate the influence of the reward model, we additionally compare against a variant that samples by directly selecting the tokens with the highest and lowest logits among the top-$k$ candidates.
Table 3: Results of training datasets constructed by different reward models.
@@ 4.5 Analysis of Reward Model
+ Qwen-2.5-Math-PRM 84.1 40.7
+ Logits 83.8 39.5
[…]
Qwen2-7B-SFT 87.9 54.2
@@
+ Qwen-2.5-Math-PRM 88.3 55.0
+ Logits 88.9 54.4
[…]
The comparison results are presented in Table  3. When using Math-Shepherd or Qwen-2.5-Math-PRM as the reward model in CGPO, the trained models still outperform the base model. Without any reward model, training on reasoning paths sampled from the top-k tokens by logit scores still yields a modest gain over the base model. These results indicate that CGPO does not depend on a stronger reward model and that its improvements primarily come from exploring and optimizing non-human-like reasoning paths. The largest gains are observed with ASPRM, likely because its token-level discrimination supports more reliable selection of non-human-like paths.
4.6 Analysis of Division Strategy
We use confidence scores to reconstruct the initial reasoning step, as well as the chosen and rejected data in the original Step-DPO-10k dataset, in order to compare human-like and non-human-like step divisions. Specifically, we truncate the initial reasoning step in the Step-DPO-10k dataset at the lowest confidence point, as in the dataset construction of CGPO. For the chosen and rejected data, we apply a stopping threshold at the lowest 2% confidence.
Model GSM8K MATH
MetaMath-Llama-8B 81.8 39.3
+ Step-DPO 84.8 40.2
+ Step-DPO-Confidence 84.4 40.2
[…]
Qwen2-7B-SFT 87.9 54.2
+ Step-DPO 88.4 54.6
+ Step-DPO-Confidence 88.1 54.5
[…]
The final results are reported in Table 4. We find that even using $\pi_{\theta}$ to split the reasoning path on outputs generated by the strong model for training, the resulting performance still surpasses the baseline. However, it does not outperform training on the $\pi_{\theta}$ own generated and split data. This result suggests that the split of its own generated outputs by the weak model is more aligned with the model, thereby leading to improved training performance.
4.7 Analysis of Stopping Threshold
We report the performance of the model under different confidence thresholds at which the sampling stops in Figure 4. For MetaMath-Llama-8B, performance keeps improving as the threshold increases. For Qwen2-7B-SFT, overall performance also shows an upward trend as the threshold rises, but the peak is achieved at a confidence threshold of 8%.
@@ 4.7 Analysis of Stopping Threshold
When we increase the confidence threshold, more samples will stop in the intermediate process rather than the generation of the EOS tag. This increases the proportion of non-human-like data, thereby enhancing the model’s exploration and learning of non-human-like reasoning paths, which ultimately improves performance. However, increasing the confidence threshold also introduces shorter lengths of chosen and rejected data, as shown in Figure 5. We argue that this is the reason for the performance decline of Qwen2-7B-SFT when the confidence threshold is greater. These results further demonstrate the necessity of optimizing the model by the intermediate reasoning path rather than by the entire process.